1. Locate the centroid of the shaded area. y = a cos TX L ㅏ늘 L 2 늘 →
Answer #1

solution:

4 a Equation of the curve, y = acos in Consider an elemental strip, which is at a distance of x units from the y-axis and who

1+cos 1 = dr - 2 2πχ sin L 27 L a 4 = x+ Jo szda - dA y-co-ordinate of the Centroid, y = a L Zaz 7 ra 8 πα 8 ye

please give me thumb up